# Profinite rigidity witnessed by Dehn fillings of cusped hyperbolic 3-manifolds

### v2

@@ 3.1. Peripheral $\widehat{\mathbb{Z}}^{\times}$-regularity
The key concept introduced in [57] is the peripheral $\widehat{\mathbb{Z}}^{\times}$-regularity, which eventually implies that a profinite isomorphism between two finite-volume hyperbolic 3-manifolds carries profinite isomorphisms between their Dehn fillings.
Recall that when $M$ is a compact, orientable 3-manifold with incompressible toral boundary, we have established a canonical isomorphism $\overline{\pi_{1}\partial_{i}M}\cong\widehat{\pi_{1}\partial_{i}M}\cong\widehat{\mathbb{Z}}\otimes_{\mathbb{Z}}\pi_{1}\partial_{i}M$ according to Proposition 2.3, where $\overline{\pi_{1}\partial_{i}M}$ denotes the closure of the peripheral subgroup in $\widehat{\pi_{1}M}$.
@@
Let $M$ and $N$ be compact, orientable 3-manifolds with incompressible toral boundary. Let $\partial_{i}M$ be a boundary component of $M$ and $\pi_{1}\partial_{i}M$ be a conjugacy representative of the corresponding peripheral subgroup. Suppose $f:\widehat{\pi_{1}M}\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}\widehat{\pi_{1}N}$ is an isomorphism of profinite groups. We say that $f$ is peripheral $\widehat{\mathbb{Z}}^{\times}$-regular at $\partial_{i}M$ if the following holds.
@@
    There exists a boundary component $\partial_{j}N$ of $N$, a conjugacy representative of the peripheral subgroup $\pi_{1}\partial_{j}N$, and an element $g\in\widehat{\pi_{1}N}$ such that $f(\overline{\pi_{1}\partial_{i}M})=g^{-1}\overline{\pi_{1}\partial_{j}N}g$.
@@
    Let $C_{g}$ denote the conjugation by $g$, and $C_{g}\circ f$ restricts to an isomorphism
[…]
    There exist an element $\lambda\in\widehat{\mathbb{Z}}^{\times}$ and an isomorphism $\psi:\pi_{1}\partial_{i}M\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}\pi_{1}\partial_{j}N$ such that $C_{g}\circ f=\lambda\otimes\psi$, where $\lambda$ denotes the scalar multiplication by $\lambda$.
One essential result in [57] is that any profinite isomorphism between finite-volume hyperbolic 3-manifolds is always peripheral $\widehat{\mathbb{Z}}^{\times}$-regular.
Theorem 3.2 ([57, Theorem 7.1] ).
Let $M$ and $N$ be orientable finite-volume hyperbolic 3-manifolds with cusps, and $f:\widehat{\pi_{1}M}\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}\widehat{\pi_{1}N}$ be an isomorphism.
  1. (1)
    $f$ respects the peripheral structure, i.e. there is a one-to-one correspondence between the boundary components of $M$ and $N$, which we simply denote as $\partial_{i}M\leftrightarrow\partial_{i}N$, such that $f(\overline{\pi_{1}\partial_{i}M})$ is a conjugate of $\overline{\pi_{1}\partial_{i}N}$ in $\widehat{\pi_{1}N}$.
  2. (2)
    $f$ is peripheral $\widehat{\mathbb{Z}}^{\times}$-regular at every boundary component of $M$:
    $C_{g_{i}}\circ f|_{\overline{\pi_{1}\partial_{i}M}}:\widehat{\mathbb{Z}}\otimes_{\mathbb{Z}}\pi_{1}\partial_{i}M\xrightarrow{\lambda_{i}\otimes\psi_{i}}\widehat{\mathbb{Z}}\otimes_{\mathbb{Z}}\pi_{1}\partial_{i}N$
    for some $g_{i}\in\widehat{\pi_{1}N}$, $\lambda_{i}\in\widehat{\mathbb{Z}}^{\times}$ and some isomorphism $\psi_{i}:\pi_{1}\partial_{i}M\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}\pi_{1}\partial_{i}N$.
In fact, we can prove that the coefficients $\lambda_{i}$ corresponding to different boundary components are equal up to a $\pm$-sign. The proof of this theorem was actually based on Liu’s observation of homological $\widehat{\mathbb{Z}}^{\times}$-regularity [35].
@@
    $N$ is also a mixed 3-manifold; in particular, $N$ is irreducible and boundary-incompressible.
@@ 3.2. Dehn filling
$\text{slp}(T^{2})=\left.\left\{\alpha\in\pi_{1}(T^{2})\left|\,\begin{gathered}\alpha\text{ is a non-zero primitive element in}\\ \text{the free abelian group }\pi_{1}(T^{2})\cong\mathbb{Z}^{2}\end{gathered}\right.\right\}\right/\alpha\sim-\alpha.$
For a non-zero primitive element $\alpha\in\pi_{1}(T^{2})$, we denote by $[\alpha]$ its image in $\text{slp}(T^{2})$. Usually, we simply denote an element in $\text{slp}(T^{2})$ by a lower-case letter, for example $c\in\text{slp}(T^{2})$.
@@
${\widehat{H}}$${\widehat{\pi_{1}M}}$${\widehat{\pi_{1}M_{(c_{i})}}}$${1.}$$\scriptstyle{\widehat{\iota}}$
Note that the image of $\widehat{\iota}:\widehat{H}\to\widehat{\pi_{1}M}$ is exactly the closure of $H$ in $\widehat{\pi_{1}M}$. Therefore, this implies (2). ∎
As a corollary of Theorem 3.2, we obtain the key technique Theorem A which is rephrased as follows. This enables us to derive information of Dehn filling from the finite quotient groups of a cusped hyperbolic 3-manfold.
@@
Suppose that $M$ and $N$ are orientable finite-volume hyperbolic 3-manifolds with cusps and that $\widehat{\pi_{1}M}\cong\widehat{\pi_{1}N}$. Then, there is a one-to-one correspondence between the cusps of $M$ and $N$ denoted by $\partial_{i}M\leftrightarrow\partial_{i}N$ ($1\leq i\leq n$), and there exist isomorphisms $\psi_{i}:\pi_{1}\partial_{i}M\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}\pi_{1}\partial_{i}N$ between the conjugacy representatives of the peripheral subgroups such that for any slopes $c_{i}\in\text{slp}^{\ast}(\partial_{i}M)$ ($1\leq i\leq n$),
@@
We choose conjugacy representatives of the peripheral subgroups $\pi_{1}\partial_{i}M\subseteq\pi_{1}M$, $\pi_{1}\partial_{j}N\subseteq\pi_{1}N$, and fix a profinite isomorphism $f:\widehat{\pi_{1}M}\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}\widehat{\pi_{1}N}$.
According to Theorem 3.2 (1), $f$ respects the peripheral structure, so up to a reordering of boundary components, $f(\overline{\pi_{1}\partial_{i}M})$ is a conjugate of $\overline{\pi_{1}\partial_{i}N}$. In addition, Theorem 3.2 (2) further implies that up to composing with a conjugation, there exist $\lambda_{i}\in\widehat{\mathbb{Z}}^{\times}$ and $\psi_{i}:\pi_{1}\partial_{i}M\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}\pi_{1}\partial_{i}N$ such that $conj\circ f|_{\overline{\pi_{1}\partial_{i}M}}=\lambda_{i}\otimes\psi_{i}$.
Let $\gamma_{i}\in\pi_{1}\partial_{i}M$ denote the representative element of $c_{i}$, and $\gamma_{i}=1$ if $c_{i}=\varnothing$. Then according to Lemma 3.5, $\widehat{\pi_{1}M_{(c_{i})}}\cong{\widehat{\pi_{1}M}}/{\overline{\langle\langle\gamma_{1},\cdots,\gamma_{n}\rangle\rangle}}$ and $\widehat{\pi_{1}N_{(\overline{\psi}_{i}(c_{i}))}}\cong\widehat{\pi_{1}N}/\overline{\langle\langle\psi_{1}(\gamma_{1}),\cdots,\psi_{n}(\gamma_{n})\rangle\rangle}$.
The peripheral $\widehat{\mathbb{Z}}^{\times}$-regularity implies that $f(\gamma_{i})$ is a conjugate of $\psi_{i}(\gamma_{i})^{\lambda_{i}}$ in $\widehat{\pi_{1}N}$, no matter $c_{i}\in\text{slp}(\partial_{i}M)$ or $c_{i}=\varnothing$. Since $\lambda_{i}\in\widehat{\mathbb{Z}}^{\times}$, $\overline{\langle\langle f(\gamma_{1}),\cdots,f(\gamma_{n})\rangle\rangle}=\overline{\langle\langle\psi_{1}(\gamma_{1})^{\lambda_{1}},\cdots,\psi_{n}(\gamma_{n})^{\lambda_{n}}\rangle\rangle}=\overline{\langle\langle\psi_{1}(\gamma_{1}),\cdots,\psi_{n}(\gamma_{n})\rangle\rangle}$ yield the same closed normal subgroup in $\widehat{\pi_{1}N}$. Therefore, $f$ descends to an isomorphism
@@
Now we prove that $f_{\star}$ respects the peripheral structure. Indeed, $\partial M_{(c_{i})}$ and $\partial N_{(\overline{\psi}(c_{i}))}$ consists of the boundary components $\partial_{j}M$ and $\partial_{j}N$, where $c_{j}=\varnothing$. The corresponding peripheral subgroups are the images of $\iota_{j_{\star}}:\pi_{1}\partial_{j}M\subseteq\pi_{1}M\to\pi_{1}M/\langle\langle\gamma_{1},\cdots,\gamma_{n}\rangle\rangle\cong\pi_{1}M_{(c_{i})}$ and $\rho_{j_{\ast}}:\pi_{1}\partial_{j}N\subseteq\pi_{1}N\to\pi_{1}N/\langle\langle\psi_{1}(\gamma_{1}),\cdots,\psi_{n}(\gamma_{n})\rangle\rangle\cong\pi_{1}N_{(\overline{\psi}(c_{i}))}$ respectively. Recall that $f$ respects the peripheral structure, i.e. $f(\overline{\pi_{1}\partial_{j}M})=g_{j}\cdot\overline{\pi_{1}\partial_{j}N}\cdot g_{j}^{-1}$ for some $g_{j}\in\widehat{\pi_{1}N}$. Desending to $f_{\star}:\widehat{\pi_{1}M_{(c_{i})}}\cong\widehat{\pi_{1}N_{(\overline{\psi}_{i}(c_{i}))}}$, we obtain that $f_{\star}(\overline{\iota_{j_{\star}}(\pi_{1}\partial_{j}M)})=\Phi(g_{j})\cdot\overline{\rho_{j_{\star}}(\pi_{1}\partial_{j}N)}\cdot\Phi(g_{j})^{-1}$, where $\Phi:\widehat{\pi_{1}N}\to\widehat{\pi_{1}N}/\overline{\langle\langle\psi_{1}(\gamma_{1}),\cdots,\psi_{n}(\gamma_{n})\rangle\rangle}\cong\widehat{\pi_{1}N_{(\overline{\psi}(c_{i}))}}$ denotes the quotient homomorphism. Thus, $f_{\star}$ also respects the peripheral structure. ∎
@@ 4.1. Detection of knot and link complements
As a special application of Theorem A, let us consider a typical example of cusped finite-volume hyperbolic 3-manifold, namely the complement of a hyperbolic link in $S^{3}$. In the following context, we also include knots when we refer to links.
@@
The same procedure also holds for a link complement which is a mixed manifold, whose boundary components belong to hyperbolic pieces. For brevity of notion, we only describe this statement for hyperbolic-type satellite knots.
@@
Let $c$ be the meridian of $K$, so that $M_{c}\cong S^{3}$. Then $\widehat{\pi_{1}M_{c}}\cong\widehat{\pi_{1}N_{\overline{\psi}(c)}}$ is the trivial group, and the residual finiteness of $\pi_{1}N_{\overline{\psi}(c)}$ (Proposition 2.1) implies that $\pi_{1}N_{\overline{\psi}(c)}$ is also the trivial group. Therefore, $N_{\overline{\psi}(c)}$ is closed, since any compact orientable 3-manifold with non-empty (non-spherical) boundary has positive first Betti number. In particular, $\partial_{0}N$ is the unique boundary component of $N$. The validity of the Poincaré conjecture then implies that $N_{\overline{\psi}(c)}\cong S^{3}$.

### v3

@@ 3. Peripheral $\widehat{\mathbb{Z}}^{\times}$-regularity and correspondence of Dehn fillings
The key concept introduced in [56] is the so-called peripheral $\widehat{\mathbb{Z}}^{\times}$-regularity. In this section, we will show that peripheral $\widehat{\mathbb{Z}}^{\times}$-regularity eventually implies that a profinite isomorphism between two cusped finite-volume hyperbolic 3-manifolds carries profinite isomorphisms between their Dehn fillings.
@@ 3.1. Peripheral $\widehat{\mathbb{Z}}^{\times}$-regularity
Recall that for a compact orientable 3-manifold $M$ with toral boundary, we have $\pi_{1}M$ sitting inside $\widehat{\pi_{1}M}$ as a subgroup via the canonical homomorphism according to Proposition 2.2, and so any peripheral subgroup $\pi_{1}\partial_{i}M$ is also naturally viewed as a subgroup of $\widehat{\pi_{1}M}$.
@@
Let $M$ and $N$ be compact, orientable 3-manifolds with incompressible toral boundary. Suppose $f:\widehat{\pi_{1}M}\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}\widehat{\pi_{1}N}$ is an isomorphism. Let $\partial_{i}M$ be a boundary component of $M$, and let $\pi_{1}\partial_{i}M$ be a conjugacy representative of the corresponding peripheral subgroup. We say that $f$ is peripheral $\widehat{\mathbb{Z}}^{\times}$-regular at $\partial_{i}M$ if the following holds.
@@
    There exist a boundary component $\partial_{j}N$ of $N$, a conjugacy representative of the peripheral subgroup $\pi_{1}\partial_{j}N$, and an element $g\in\widehat{\pi_{1}N}$ such that $f(\overline{\pi_{1}\partial_{i}M})=g^{-1}\cdot\overline{\pi_{1}\partial_{j}N}\cdot g$.
@@
    There exist an element $\lambda\in\widehat{\mathbb{Z}}^{\times}$ and an isomorphism $\psi:\pi_{1}\partial_{i}M\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}\pi_{1}\partial_{j}N$ such that $f(\gamma)=g^{-1}\cdot\psi(\gamma)^{\lambda}\cdot g$ for any $\gamma\in\pi_{1}\partial_{i}M$.
One essential result in [56] is that any profinite isomorphism between cusped finite-volume hyperbolic 3-manifolds is always peripheral $\widehat{\mathbb{Z}}^{\times}$-regular.
Theorem 3.2 ([56, Theorem 7.2]).
Let $M$ and $N$ be orientable cusped finite-volume hyperbolic 3-manifolds, and suppose $f:\widehat{\pi_{1}M}\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}\widehat{\pi_{1}N}$ is an isomorphism. Then, $f$ respects the peripheral structure and $f$ is peripheral $\widehat{\mathbb{Z}}^{\times}$-regular at all boundary components of $M$.
Note that in the setting of Theorem 3.2, given a boundary component $\partial_{i}M$, the corresponding boundary component $\partial_{j}N$ appearing in (1) of Definition 3.1 is unique by [53, Lemma 4.5], which gives a one-to-one correspondence between the boundary components of $M$ and $N$. The proof of this theorem was actually based on Liu’s observation of homological $\widehat{\mathbb{Z}}^{\times}$-regularity [35].
@@
Let $M$ be a mixed 3-manifold, and let $\partial_{1}^{hyp}M,\cdots,\partial_{n}^{hyp}M$ be the boundary components of $M$ belonging to hyperbolic JSJ-pieces. Suppose $N$ is a compact, orientable 3-manifold, and $f:\widehat{\pi_{1}M}\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}\widehat{\pi_{1}N}$ is an isomorphism.
@@
    NN is also a mixed 3-manifold; in particular, $N$ is irreducible and boundary-incompressible.
@@ 3.2. Dehn filling
[…]
For a non-zero primitive element $\alpha\in\pi_{1}(T^{2})$, we denote by $\alpha]$ its image in $\text{slp}(T^{2})$. Usually, we simply denote an element in $\text{slp}(T^{2})$ by a lower-case letter, for example $c\in\text{slp}(T^{2})$.
@@
[…]
Note that the image of $\widehat{\iota}:\widehat{H}\to\widehat{\pi_{1}M}$ is exactly the closure of $H$ in $\widehat{\pi_{1}M}$, which is the closed normal subgroup generated by $\gamma_{1},\cdots,\gamma_{n}$, so this implies (2). ∎
As an application of Theorem 3.2, we obtain the key technique Theorem A which is rephrased as follows.
@@
Suppose that $M$ and $N$ are orientable cusped finite-volume hyperbolic 3-manifolds and that $\widehat{\pi_{1}M}\cong\widehat{\pi_{1}N}$. Then, there is a one-to-one correspondence between the cusps of $M$ and $N$ denoted by $\partial_{i}M\leftrightarrow\partial_{i}N$ ($1\leq i\leq n$) up to a reordering, and there exist isomorphisms $\psi_{i}:\pi_{1}\partial_{i}M\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}\pi_{1}\partial_{i}N$ between the conjugacy representatives of the peripheral subgroups such that for any slopes $c_{i}\in\text{slp}^{\ast}(\partial_{i}M)$ ($1\leq i\leq n$),
@@
We choose conjugacy representatives of the peripheral subgroups $\pi_{1}\partial_{i}M\subseteq\pi_{1}M$, $\pi_{1}\partial_{j}N\subseteq\pi_{1}N$, and fix a profinite isomorphism $f:\widehat{\pi_{1}M}\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}\widehat{\pi_{1}N}$. According to Theorem 3.2, up to a reordering, $f(\overline{\pi_{1}\partial_{i}M})$ is a conjugate of $\overline{\pi_{1}\partial_{i}N}$ for each $i$, and there exist coefficients $\lambda_{i}\in\widehat{\mathbb{Z}}^{\times}$ and isomorphisms $\psi_{i}:\pi_{1}\partial_{i}M\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}\pi_{1}\partial_{i}N$ such that for any $\gamma_{i}\in\pi_{1}\partial_{i}M$, $f(\gamma_{i})$ is a conjugate of $\psi_{i}(\gamma_{i})^{\lambda_{i}}$ in $\widehat{\pi_{1}N}$. We show that the isomorphisms $\psi_{i}$ satisfy the requirement of the theorem.
Let $\gamma_{i}\in\pi_{1}\partial_{i}M$ denote a representative element of $c_{i}$, and $\gamma_{i}=1$ if $c_{i}=\varnothing$. Since each $\lambda_{i}\in\widehat{\mathbb{Z}}^{\times}$, by the reasoning in § 2.2, we have $f(\overline{\langle\!\langle\gamma_{1},\cdots,\gamma_{n}\rangle\!\rangle})=\overline{\langle\!\langle f(\gamma_{1}),\cdots,f(\gamma_{n})\rangle\!\rangle}=\overline{\langle\!\langle\psi_{1}(\gamma_{1})^{\lambda_{1}},\cdots,\psi_{n}(\gamma_{n})^{\lambda_{n}}\rangle\!\rangle}=\overline{\langle\!\langle\psi_{1}(\gamma_{1}),\cdots,\psi_{n}(\gamma_{n})\rangle\!\rangle}$. Moreover, according to Lemma 3.5, $\widehat{\pi_{1}M_{(c_{i})}}\cong{\widehat{\pi_{1}M}}/{\overline{\langle\!\langle\gamma_{1},\cdots,\gamma_{n}\rangle\!\rangle}}$ and $\widehat{\pi_{1}N_{(\overline{\psi}_{i}(c_{i}))}}\cong\widehat{\pi_{1}N}/\overline{\langle\!\langle\psi_{1}(\gamma_{1}),\cdots,\psi_{n}(\gamma_{n})\rangle\!\rangle}$. Therefore, $f$ descends to an isomorphism
@@
Now we prove that $f_{\star}$ respects the peripheral structure. Indeed, $\partial M_{(c_{i})}$ and $\partial N_{(\overline{\psi}(c_{i}))}$ consists of the boundary components $\partial_{j}M$ and $\partial_{j}N$, where $c_{j}=\varnothing$. The corresponding peripheral subgroups are the images of $\iota_{j_{\star}}:\pi_{1}\partial_{j}M\subseteq\pi_{1}M\to\pi_{1}M/\langle\!\langle\gamma_{1},\cdots,\gamma_{n}\rangle\!\rangle\cong\pi_{1}M_{(c_{i})}$ and $\rho_{j_{\ast}}:\pi_{1}\partial_{j}N\subseteq\pi_{1}N\to\pi_{1}N/\langle\!\langle\psi_{1}(\gamma_{1}),\cdots,\psi_{n}(\gamma_{n})\rangle\!\rangle\cong\pi_{1}N_{(\overline{\psi}(c_{i}))}$ respectively. Recall that $f$ respects the peripheral structure, i.e. $f(\overline{\pi_{1}\partial_{j}M})=g_{j}\cdot\overline{\pi_{1}\partial_{j}N}\cdot g_{j}^{-1}$ for some $g_{j}\in\widehat{\pi_{1}N}$. Desending to $f_{\star}:\widehat{\pi_{1}M_{(c_{i})}}\cong\widehat{\pi_{1}N_{(\overline{\psi}_{i}(c_{i}))}}$, we have $f_{\star}(\overline{\iota_{j_{\star}}(\pi_{1}\partial_{j}M)})=\phi(g_{j})\cdot\overline{\rho_{j_{\star}}(\pi_{1}\partial_{j}N)}\cdot\phi(g_{j})^{-1}$, where $\phi:\widehat{\pi_{1}N}\to\widehat{\pi_{1}N}/\overline{\langle\!\langle\psi_{1}(\gamma_{1}),\cdots,\psi_{n}(\gamma_{n})\rangle\!\rangle}\cong\widehat{\pi_{1}N_{(\overline{\psi}(c_{i}))}}$ denotes the quotient homomorphism. Thus, $f_{\star}$ also respects the peripheral structure. ∎
@@ 4.1. Detection of knot and link complements
As a special application of Theorem A, let us consider a typical example of cusped finite-volume hyperbolic 3-manifold, namely the complement of a hyperbolic link in $S^{3}$. In the following context, all links are tame, and we also include knots when we refer to links.
@@
The same procedure also holds for a link complement which is a mixed manifold, whose boundary components belong to hyperbolic pieces. For brevity of notation, we only describe this statement for hyperbolic-type satellite knots.
@@
Let $c$ be the meridian of $K$, so that $M_{c}\cong S^{3}$. Then $\widehat{\pi_{1}M_{c}}\cong\widehat{\pi_{1}N_{\overline{\psi}(c)}}$ is the trivial group, and the residual finiteness of $\pi_{1}N_{\overline{\psi}(c)}$ (Proposition 2.2) implies that $\pi_{1}N_{\overline{\psi}(c)}$ is also the trivial group. Therefore, $N_{\overline{\psi}(c)}$ is closed, since any compact orientable 3-manifold with non-empty non-spherical boundary has positive first Betti number. In particular, $\partial_{0}N$ is the unique boundary component of $N$. Again, the validity of the Poincaré conjecture implies that $N_{\overline{\psi}(c)}\cong S^{3}$.